# DSSmoothing: Toward Certified Dataset Ownership Verification for Pre-trained Language Models via Dual-Space Smoothing

### v1

@@ Appendix A Proof of Theorem 4
Assume that the permutation noise $\rho\sim\mathcal{U}[-\lambda,\lambda]$. Denote by $W\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)$ the watermark robustness (WR) of the suspicious model $f_{\bm{\theta}}$ under permutation-space smoothing, as defined in Eq. (7), which is $L$-Lipschitz continuous under the $\|\cdot\|$ norm. Let $\delta_{p}$ represent the permutation-space watermark perturbations, and define $r_{p}=\max_{k=1,\ldots,K}\|\delta_{p}\|_{1}$ as the maximum perturbation magnitude in this space. Then, text dataset ownership verification in the permutation space is guaranteed if
@@
(13) $W\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)>P_{C}^{(J-m-\lfloor\alpha_{0}(J-m+1)\rfloor)}(\mathcal{P}_{\rho}).$
[…]
(14) $\big|W\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)-\hat{W}\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)\big|\leq L\cdot r_{p}\;\Longrightarrow\;W\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)\geq\hat{W}\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)-L\cdot r_{p}.$
To ensure the condition in Eq. (13), it is sufficient to require
(15) $\hat{W}\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)-L\cdot r_{p}>P_{C}^{(J-m-\lfloor\alpha_{0}(J-m+1)\rfloor)}(\mathcal{P}_{\rho}),$
which corresponds exactly to the robustness condition in Eq. (12). Therefore, dataset ownership verification in the permutation space is guaranteed under the stated robustness bound. ∎
@@
Proof of Theorem 4. According to Lemma 1, the uniform-based smoothed classifier $\tilde{h}(\bm{U},\bm{W})$ is $\tfrac{1}{2\lambda}$-Lipschitz in $\bm{U}$ under the $\ell_{1}$ norm. Combining this lemma with Lemma 2, and substituting $L=\tfrac{1}{2\lambda}$ into Eq. (12), we obtain Eq. (16) in Lemma 3. Together with the embedding-space result established in CertDW (Qiao et al., 2025), this completes the proof of Theorem 4. ∎
@@ B.2. Randomized Smoothing
We here provide the formal definition of randomized smoothing (RS) (Cohen et al., 2019), a certified defense technique that leverages Gaussian perturbations to provide provable robustness guarantees. Specifically, for any input $\bm{x}$, a smoothed classifier $g$ is constructed by adding Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(0,\sigma^{2}I)$, where $\sigma$ is a hyperparameter controlling the robustness level. This classifier outputs the class most frequently predicted by the base model $f$ under noise perturbations: $g(\bm{x}):=\arg\max_{y\in\mathcal{Y}}\mathbb{P}(f(\bm{x}+\bm{\epsilon})=y)$. Based on the Neyman-Pearson Lemma (Neyman and Pearson, 1933), Cohen et al. (Cohen et al., 2019) established the certification mechanism of this method under $\ell_{2}$ perturbations. Let $y_{A}$ denote the most probable class under noise conditions ($i.e.$, $y_{A}=\arg\max_{y}\mathbb{P}(f(\bm{x}+\bm{\epsilon})=y)$), with corresponding probability denoted as $P_{A}=\mathbb{P}(f(\bm{x}+\bm{\epsilon})=y_{A})$; the maximum probability of the runner-up class is denoted as $P_{B}=\max_{y\neq y_{A}}\mathbb{P}(f(\bm{x}+\bm{\epsilon})=y)$. Theoretical analysis shows that when an adversarial perturbation $\bm{\tau}$ satisfies $\|\bm{\tau}\|_{2}<r$, the classification result remains stable ($i.e.$, $g(\bm{x}+\bm{\tau})=y_{A}$), where the certified radius $r$ is computed as:
(17) $r:=\frac{\sigma}{2}\left(\Phi^{-1}(P_{A})-\Phi^{-1}(P_{B})\right),$
@@ Appendix E The Analysis of Computational Complexity
Complexity of Watermarked Dataset Generation. Let $N$ denote the size of the text dataset and $\gamma$ be the poisoning rate. In the generation stage, we only need to embed watermarks into $\gamma\cdot N$ samples, resulting in a complexity of $\mathcal{O}(\gamma\cdot N)$. In addition, our method introduces perturbations in both the embedding space and the permutation space: in the embedding space, we optimize trigger embeddings through local pooling and iterative scaling, with a complexity of $\mathcal{O}(|P|\cdot u\cdot d)$, where $|P|$ is the number of trigger positions, $u$ is the window size, and $d$ is the embedding dimension; in the permutation space, we perform group-based reordering over a sequence of length $n$, with a complexity of $\mathcal{O}(n)$. Since these additional costs are negligible compared with the overall dataset scale, the total complexity can still be regarded as $\mathcal{O}(\gamma\cdot N)$, and the extra time overhead is almost negligible in large-scale training pipelines.
Complexity of Ownership Verification. Let $J$, $W_{a}$, and $I_{n}$ denote the numbers of benign models, watermarked models, and independent models, respectively. Training these models requires complexities of $\mathcal{O}(J)$, $\mathcal{O}(W_{a})$, and $\mathcal{O}(I_{n})$, respectively, and this process can be significantly accelerated through parallel computation. In the verification stage, we need to compute both the watermark robustness (WR) and the principal probability (PP) under dual-space randomized smoothing. For the suspicious model, WR computation involves performing Monte Carlo sampling on watermarked samples from each class while simultaneously injecting noise into both embedding and permutation spaces, which results in a complexity of $\mathcal{O}(K\cdot M\cdot\mathcal{C}(f))$, where $K$ is the number of classes, $M$ is the number of samples, and $\mathcal{C}(f)$ denotes the cost of a single forward pass. For benign models, PP computation also relies on dual-space smoothing and requires estimating prediction distributions across multiple models, with a complexity of $\mathcal{O}(J\cdot K\cdot M\cdot\mathcal{C}(g))$. In practice, both WR and PP calculations only involve forward inference and can be highly parallelized on GPUs, making their cost negligible compared with model training. For example, on common text classification datasets, training a benign or watermarked model typically takes several minutes, whereas computing WR and PP under dual-space smoothing can be completed within a few seconds, thus incurring almost no additional overhead.
Appendix F Additional Results on the AG News Dataset
@@
Appendix G Additional Effectiveness Results on the AG News Dataset
@@ Appendix H Effects of the Number of Benign Models
Appendix I Model-level Transferability on AG Dataset
@@
Appendix J Societal Impacts
This paper focuses on protecting public datasets used in pre-trained language model (PLM) training. We identify that existing dataset ownership verification (DOV) methods are vulnerable to perturbations and propose a robust, certifiable approach that enables privacy-preserving and traceable ownership verification. The method is purely defensive, aiming to promote fair data usage, trustworthy sharing, and responsible AI development. Although it employs a trigger-based watermarking mechanism similar in form to backdoor attacks, its use is strictly limited to legitimate and authorized verification scenarios. When applied in controlled and compliant environments, it introduces no additional security or privacy risks, thereby contributing positively to trustworthy data governance and the sustainable development of transparent AI systems.

### v2

@@ Appendix A Proof of Theorem 4
Assume that the permutation noise $\rho\sim\mathcal{U}[-\lambda,\lambda]$. Denote by $W\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)$ the watermark robustness (WR) of the suspicious model $f_{\bm{\theta}}$ under permutation-space smoothing, as defined in Eq. (7), which is $L$-Lipschitz continuous under the $\|\cdot\|_{1}$ norm. Let $\delta_{p}$ represent the permutation-space watermark perturbations, and define $r_{p}=\max_{k=1,\ldots,K}\|\delta_{p}\|_{1}$ as the maximum perturbation magnitude in this space. Then, text dataset ownership verification in the permutation space is guaranteed if
@@
(13) $\hat{W}\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)>P_{C}^{(J-m-\lfloor\alpha_{0}(J-m+1)\rfloor)}(g_{w},\mathcal{P}_{\rho}).$
Since the WR functional is $L$-Lipschitz continuous w.r.t. the input under the $\|\cdot\|_{1}$ norm and $\|\widehat{\bm{U}}_{k}-{\bm{U}}_{k}\|_{1}\leq r_{p}$, we have
(14) $\hat{W}\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)\geq W\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)-L\,r_{p}.$
Therefore, a sufficient condition for Eq. (13) is
$W\!\big(f_{\bm{\theta}},\mathcal{P}_{\rho}\big)-L\,r_{p}>P_{C}^{(J-m-\lfloor\alpha_{0}(J-m+1)\rfloor)}(g_{w},\mathcal{P}_{\rho}),$
which is equivalent to Eq. (12). This completes the proof. ∎
@@
Proof of Theorem 4. According to Lemma 1, the uniform-based smoothed classifier $\tilde{h}(\bm{U},\bm{W})$ is $\tfrac{1}{2\lambda}$-Lipschitz in $\bm{U}$ under the $\ell_{1}$ norm. Combining this Lemma 1 with Lemma 2, and substituting $L=\tfrac{1}{2\lambda}$ into Eq. (12), we obtain Eq. (15) in Lemma 3. Together with the embedding-space result established in CertDW (Qiao et al., 2025), this completes the proof of Theorem 4. ∎
@@ B.2. Randomized Smoothing
We here provide the formal definition of randomized smoothing (RS) (Cohen et al., 2019), a certified defense technique that leverages Gaussian perturbations to provide provable robustness guarantees. Specifically, for any input $\bm{x}$, a smoothed classifier $g$ is constructed by adding Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(0,\sigma^{2}I)$, where $\sigma$ is a hyperparameter controlling the robustness level. This classifier outputs the class most frequently predicted by the base model $f$ under noise perturbations: $g(\bm{x})\triangleq\arg\max_{y\in\mathcal{Y}}\mathbb{P}(f(\bm{x}+\bm{\epsilon})=y)$. Based on the Neyman-Pearson Lemma (Neyman and Pearson, 1933), Cohen et al. (Cohen et al., 2019) established the certification mechanism of this method under $\ell_{2}$ perturbations. Let $y_{A}$ denote the most probable class under noise conditions ($i.e.$, $y_{A}=\arg\max_{y}\mathbb{P}(f(\bm{x}+\bm{\epsilon})=y)$), with corresponding probability denoted as $P_{A}=\mathbb{P}(f(\bm{x}+\bm{\epsilon})=y_{A})$; the maximum probability of the runner-up class is denoted as $P_{B}=\max_{y\neq y_{A}}\mathbb{P}(f(\bm{x}+\bm{\epsilon})=y)$. Theoretical analysis shows that when an adversarial perturbation $\bm{\tau}$ satisfies $\|\bm{\tau}\|_{2}<r$, the classification result remains stable ($i.e.$, $g(\bm{x}+\bm{\tau})=y_{A}$), where the certified radius $r$ is computed as:
(16) $r\triangleq\frac{\sigma}{2}\left(\Phi^{-1}(P_{A})-\Phi^{-1}(P_{B})\right),$
@@ C.1. Certified Backdoor Robustness
Figure 11. Effects of the number of benign models.
Line plot showing how verification performance varies with the number of benign reference models.
@@ Appendix E The Analysis of Computational Complexity
Complexity of Watermarked Dataset Generation. Let $N$ denote the size of the text dataset and $\gamma$ the poisoning rate. In the generation stage, we only need to embed watermarks into $\gamma N$ samples, resulting in a complexity of $\mathcal{O}(\gamma N)$. Moreover, we introduce perturbations in both the embedding and permutation spaces: in the embedding space, we optimize trigger embeddings via local pooling and iterative scaling, with a complexity of $\mathcal{O}(|P|\cdot u\cdot d)$, where $|P|$ is the number of trigger positions, $u$ is the window size, and $d$ is the embedding dimension; in the permutation space, we perform local group-based reordering for a sequence of length $n$, which incurs a complexity of $\mathcal{O}(n)$. Since these two components impose relatively minor overhead compared to the overall data processing scale, the total complexity can still be regarded as $\mathcal{O}(\gamma N)$, with negligible additional runtime overhead.
Complexity of Ownership Verification. Let $J$, $W_{a}$, and $I_{n}$ denote the numbers of benign models, watermarked models, and independent models, respectively. Training these models incurs complexities of $\mathcal{O}(J)$, $\mathcal{O}(W_{a})$, and $\mathcal{O}(I_{n})$, and can be significantly accelerated via parallel training. During verification, we compute the watermark robustness (WR) and the principal probability (PP) under dual-space randomized smoothing. For the suspicious model, WR is estimated by running $M$ Monte Carlo trials on a watermarked example from each class while injecting noise into both the embedding and permutation spaces, resulting in a complexity of $\mathcal{O}(K\cdot M\cdot\mathcal{C}(f))$, where $K$ is the number of classes and $\mathcal{C}(f)$ denotes the cost of a single forward pass. For benign models, PP is computed by estimating prediction distributions across $J$ models, with a complexity of $\mathcal{O}(J\cdot K\cdot M\cdot\mathcal{C}(g))$. Since WR/PP only involve forward inference and are highly parallelizable on GPUs, their cost is typically negligible compared to model training. For example, on the AG dataset with BERT, training a benign model takes about 5 minutes, while computing WR and PP takes about 30 seconds, making the verification overhead relatively small and acceptable.
Appendix F Additional Results on the AG
@@
Appendix G Additional Effectiveness Results on the AG
@@ Appendix H Effects of the Number of Benign Models
Two subfigures for the AG dataset comparing watermark robustness (WR) under embedding-space perturbations. The left panel uses the BadWord trigger and the right panel uses AddSent; the bullet marks the original watermarked sample and the curves show WR under noise and adversarial directions.
@@
Two subfigures for the AG dataset comparing watermark robustness (WR) of our method under embedding-space perturbations. The left panel uses the BadWord trigger and the right panel uses AddSent; the bullet marks the original watermarked sample and the curves show WR under noise and adversarial directions.
@@
Two heatmaps for the AG dataset. The left heatmap shows verification success rate (VSR) and the right heatmap shows watermark certification accuracy (WCA), comparing different model architectures where rows correspond to benign models and columns correspond to watermarked models.
Appendix I Model-level Transferability on AG